# Why Static Friction Decreases From Single to Multi-asperity Contacts

### v1

@@ Abstract
The key parameter for describing frictional strength at the onset of sliding is the static friction coefficient. Yet, how the static friction coefficient at the macroscale emerges from contacting asperities at the microscale is still an open problem. Here, we present friction experiments in which the normal load was varied over more than three orders of magnitude, so that a transition from a single asperity contact at low loads to multi-asperity contacts at high loads was achieved. We find a remarkable drop in static friction coefficient with increasing normal load. Using a simple stick-slip transition model we identify the presence of pre-sliding and subcritical contact points as the cause of smaller static friction coefficient at increased normal loads. Our measurements and model bridge the gap between friction behavior commonly observed in atomic force microscopy (AFM) experiments at microscopic forces, and industrially relevant multi-asperity contact interfaces loaded with macroscopic forces.
@@
In this Letter, we experimentally demonstrate that the static friction coefficient decreases as the interface grows from single contact to multi-asperity contact by increasing the normal load applied to a silicon-on-silicon contact. The observed decreasing trend of static friction coefficient with increasing normal load can be captured by a simple stick-slip transition model, analogous to a fiber bundle model, based on the local normal contact stress map obtained from boundary integral contact calculations without adjustable parameters. We find that three types of contact points can be identified when the interface is subjected to the maximal tangential force it can withstand. We describe these three types of contact asperities as “critical”, “pre-sliding” and “subcritical”. While single-asperity interfaces are always critical at the onset of slip, we show that the emergence of pre-sliding and subcritical contact asperities at multi-asperity interfaces causes the static friction coefficient to decrease.
@@
Figure 1: Experimental system. (a) A 3 mm-diameter silicon ball inside a humidity-controlled chamber is clamped to the geometry of a rheometer and brought into contact with a homemade Si cantilever that is made by gluing a piece of 1×1 cm p-type silicon (100) wafer (University Wafer) to a brass sheet. The normal stiffness of the Si cantilever can be tuned by changing the width and thickness of the used brass sheet. The Si cantilever is mounted on a stage, which was attached to the weighing pan of an analytical balance by dead weight. By rotating the geometry at a constant angular velocity, $\omega$, a sliding speed of $\omega r$ ($r=10\text{ mm}$) is imposed at the Si-on-Si contact interface. The corresponding friction force ($F_{f}$) and the normal force ($N$) can be simultaneously measured by the rheometer and the balance. The ratio of friction force and normal load gives the coefficient of friction, $\mu={F_{f}}/{N}$. (b) An example of the slide-hold-slide experiments protocol with a normal load of 24 mN and a hold time of 100 s. $F_{S}$ and $F_{D}$ represent the maximum static friction force and the average dynamic friction force after the sliding is reinitiated, respectively. (c) Reproduction of the nanoscale friction coefficient evolution with imposed displacement, from previous AFM experiments [15].
@@
We study silicon-on-silicon friction using a rheometer (DSR 502, Anton Paar) in combination with an analytical balance (Quintix, Sartorius) inside a humidity-controlled chamber, as shown in Fig. 1(a). Before the friction experiments, the (naturally oxidized) silicon balls (Goodfellow) were sonicated in ethanol and then Milli-Q water, followed by nitrogen flow drying. To compare with the highly hydroxylated silicon surfaces used in previous AFM experiments [16, 15], the silicon balls and the silicon cantilevers were then treated by oxygen plasma cleaning inside a reactor (Zepto One, Diener Electronic) [23]. In the friction experiments, both the plasma-cleaned silicon ball and the plasma-cleaned silicon cantilever were initially equilibrated for one hour inside the dry chamber before contact was formed between them. The relative humidity inside the chamber was stabilized at 0.8% by continuously flowing dry nitrogen. During the sliding, the rheometer and the balance simultaneously measured the friction force and normal load at the contact interface, respectively. By using silicon cantilevers of different normal stiffness, the normal loads were precisely varied from 0.1 to 100 mN. The imposed sliding speed was fixed at 0.3 µm/s for all the measurements. A slide-hold-slide (SHS) measurement procedure was employed (see Fig. 1(b)): we imposed a stationary state with the rheometer after the initial sliding state, followed by reinitiating the sliding in the same direction. The static friction force ($F_{S}$) and dynamic friction force ($F_{D}$) when sliding was restarted, were recorded to calculate the corresponding static friction coefficient, $\mu_{S}$, the dynamic friction coefficient, $\mu_{D}$, and the normalized static friction coefficient (also called friction drop), $\mu_{S}/\mu_{D}$, which equals to $F_{S}/F_{D}$. All sliding strokes were measured on previously untouched areas of the silicon wafer and were kept short at 10 µm to minimize the influence of wear (Fig. S3).
@@
Both the static and the dynamic friction coefficients, defined as the ratio of friction force to normal force, are widely assumed to be independent of the normal load [35, 36, 37]. Surprisingly, we observe that the static friction coefficient gradually decreases with increasing normal load as shown with black data points in Fig. 2, indicating a smaller friction coefficient at higher normal loads. Notably, the single-asperity nanoscale static friction (leftmost black circle in Fig. 2), measured from previous AFM experiments (reproduced in Fig. 1(c)) [15], aligns well with our observed qualitative decreasing trend of friction drop with increasing normal load.
@@
Figure 2: The ratio of the static friction force to the dynamic friction force (also known as friction drop) as a function of normal load. The black and red data points correspond to a hold time of 100 s and 10 s, respectively. Different symbols represent the measurements using Si cantilevers of different normal stiffness and each data point represents the average of three measurements. The leftmost red and black solid circle data points, measured from AFM experiments, are extracted from Fig. S5 in Ref. [15]. The shaded area shows the friction drop calculated from our model using three different surface topographies (shown in Fig. S1) of the experimentally used Si balls, with red and gray indicating the result of a hold time of 100 s and 10 s, respectively. The vertical lines indicate the normal load corresponding to the normal stress distribution shown in Fig. 3.
@@
To provide insight into the normal load dependence of the friction drop, we propose a simple stick-slip transition model. In the model, we assume that the macroscopic friction force is the sum of all the microscopic interfacial shear forces generated by single asperity contacts. We also assume that the asperity-scale frictional behavior is characterized by an instant drop in friction upon slip, the magnitude of which matches that observed in the AFM friction experiments shown in Fig. 1(c). In other words, the local ratio of shear stress to normal stress drops at the onset of local slip, in accordance with the static and dynamic friction coefficient ($\mu^{S}$ and $\mu^{D}$, respectively) observed in the AFM experiments (Fig. 1(c)). The applied shear force is assumed to be distributed homogeneously over all the contact asperities before any asperity starts to slip. The excess friction force that is generated when a local contact asperity enters slip, is uniformly redistributed over all the contact asperities at the interface that have not undergone slip yet.
@@
To estimate the local normal stress at the interface, we conduct elastic contact calculations using an open-source implementation of the boundary integral method (BIM) [38]. The inputs for our contact calculations include the normal load, the silicon surface topography measured using AFM after friction experiments, and the mechanical properties of silicon: Young’s modulus: $E=130\pm 3\,\text{GPa}$, Poisson’s ratio: $\nu=0.26$. Our use of elastic contact is reasonable here, since there is mainly elastic deformation at the contact interface (see Fig. 3) [30, 23, 39]. Using the calculated local normal stress distribution and the assumed relations between local shear stress and local normal stress described above, the model predicts how the total friction force, which is externally imposed, is distributed over the contact points and how this distribution of local shear force evolves as the applied macroscopic tangential force is increased. As the local friction coefficient at each asperity is assumed to be uniform, it is the normal stress distribution that completely determines the frictional behavior. This enables us to predict the frictional response of the same interface at various normal loads, from which we can compute the static friction coefficient and friction drop.
@@
In other words, our model is analogous to a fiber bundle model with equal load sharing [40], where the failure of a bundle of fibers corresponds to the failure of contact asperities, with their strength governed by the local normal stress. While sliding heterogeneity can originate from elastic interaction among contact asperities [11], we assume that these interactions among asperities during the onset of the sliding process are negligible in our system. Indeed, the interfacial nanometer-sized contact deformation ($\sim$ 5 nm at a normal load of 10 mN based on BIM contact calculation as mentioned above) is considerably smaller in comparison to the micrometer-scale distance between contact asperities, which supports our assumption of equal load sharing.
@@
To quantify ball-to-ball topography variations, we use three different topographies extracted from silicon balls that were used in the experiments (see Fig. S1). For each topography and normal load, the model predicts the friction drop. The gray shaded area in Fig. 2 corresponds to the range of predicted friction drop values based on the three topographies, with the nanoscale static and dynamic friction coefficients measured at a hold time of 100 s in AFM experiments. The stick-slip transition model qualitatively reproduces the decrease of the friction drop with increasing normal load and quantitatively matches the experiments without any fitting parameter.
@@
To evaluate the robustness of our model against variations in the system characteristics, we conduct additional experiments with a hold time of 10 s instead of 100 s, and report the corresponding friction drop with red data points in Fig. 2. A decrease of the friction drop with increasing normal load is observed, similar to the experiments for a hold time of 100 s. Furthermore, it is noteworthy that the measured friction drop for a hold time of 100 s is larger than that for a hold time of 10 s under the same normal force. The larger static friction coefficient, corresponding to a greater friction drop, is expected: a longer hold time during the SHS procedure can enhance the contact strength by forming more interfacial siloxane bonds [15, 23], which results in a higher static friction coefficient. Moreover, our stick-slip transition model (red-shaded area in Fig. 2) with the nanoscale static and dynamic friction coefficients measured at a hold time of 10 s from previous AFM experiments [15] (leftmost red circle in Fig. 2, see also Fig. S2), still captures successfully the evolution of friction drop with the normal force for the experiments with 10 s hold time.
@@
One should note that there is still a gap between the minimum normal load used in the simulation and AFM experiments in Fig. 2 (marked by black and red horizontal dot lines), which will be discussed later.
@@
To further elucidate the relation between the normal load, the local normal stress distribution, and the macroscopic frictional response, we plot the probability density function, $p(\sigma)$, of the normal stress at the interface, obtained from elastic contact calculations with various normal loads in Fig. 3. With increasing normal load, the probability density function of the local normal stress evolves from being discrete (the black curve, 0.1 mN) to continuous (the yellow curve, 10mN). This directly relates to the transition from single to multi-asperity contact interfaces. At lowest normal loads, only a few contacts are formed, and the macroscopic behavior may be dominated by those asperities. At largest normal loads, however, there is a more continuous distribution of normal (and thus shear) stresses at the interface. The normal loads corresponding to the distribution shown in Fig. 3 are indicated with vertical lines of the same color in Fig. 2.
@@
Figure 4: Local stick-slip transition at the critical state of the interface from numerical computation. (a) Friction coefficient vs displacement curve for the interface at a normal load of 0.1 mN. The black circle corresponds to the critical interfacial state where no more additional tangential load can be added. (b) to (e) Distribution of local friction coefficient for a single-asperity interface (b), two-asperity interface (c)(d), and multi-asperity interface under a normal load of 10 mN (e), respectively. The brown, yellow, and green correspond to the critical, pre-sliding, and subcritical asperities with local friction coefficients of $\mu^{D}$, $\mu^{S}$, and $\mu^{0}$, respectively. $\mu^{0}$ is smaller than $\mu^{S}$. $n^{\mu}$ and $n^{tot}$ represent the number of asperities under a local friction coefficient of $\mu$ ($\mu=\mu^{S}$, $\mu^{D}$, or $\mu^{0}$) and the total number of asperities, respectively. Movies of the evolution of the distribution of local friction coefficient with increasing imposed displacement are available online (f) The state of the contacting asperities at the critical interfacial state for a normal load of 10 mN. (g) The fraction of the total normal force (solid lines) and the total friction force (dashed lines) generated by different types of contact asperities in the interfaces under various normal loads. $N^{\mu}$ and $N$ represent the shared normal load on those contact asperities under a local friction coefficient of $\mu$ ($\mu=\mu^{S}$, $\mu^{D}$, or $\mu^{0}$) and the total applied normal load, respectively. $F_{f}^{\mu}$ and $F_{f}$ represent the same quantity but for the shear load.
@@
To better understand how the differences in normal stress distribution impact the onset of sliding, we focus on the critical state of the interface, where the peak macroscopic friction coefficient is reached under an imposed tangential force corresponding to an imposed lateral displacement of $\delta_{\text{peak}}$, see a typical friction versus displacement curve from a numerical simulation in Fig. 4(a). We plot the distribution of local friction coefficients across all the contact asperities at the critical state of the interface in Fig. 4(b)-(e). For a single asperity contact (Fig. 4(b)), the asperity will reach the local maximum static friction coefficient, $\mu^{S}$, resulting in an interfacial friction coefficient equal to the local friction coefficient ($\mu=\mu^{S}$). In this limiting case of a single asperity contact, one recovers the frictional behavior as measured from AFM. In our model, this means that below a normal load for which the real contact area is of the scale of the input AFM topography resolution (21 nm), the macroscopic friction coefficient is not dependent on the normal load. Such an extreme case is marked by black and red horizontal dot lines in Fig. 2.
@@
The situation becomes more complex when considering multi-asperity contact interfaces, as each asperity potentially experiences different local normal forces (and thus different local friction coefficients at the critical interfacial state). We stress that, within our model, any distribution of local friction coefficients (between 00 and $\mu^{S}$) at a multi-contact interface can only result in a decrease in the macroscopic static friction coefficient as compared to the maximum static friction coefficient ($\mu^{S}$) at single-asperity level (Fig. 4(b)). Hereafter we investigate the details of the critical interfacial state for multi-asperity interfaces and its influence on the macroscopic friction coefficient, starting with two contacting asperities.
@@
For an interface composed of two contacting asperities, each carrying a different normal force, two distinct scenarios arise at the critical interfacial state corresponding to maximal static friction force. In the first scenario, the normal force carried by each asperity varies significantly (Fig. 4(c)) such that the weaker asperity, carrying a smaller normal force, has already slipped and experiences local friction coefficient $\mu^{D}$, while the stronger asperity with the higher normal force remains stuck with a local friction coefficient of $\mu^{S}$. We can then calculate the interfacial friction coefficient as $\mu=(\mu^{D}N^{D}+\mu^{S}N^{S})/N$, where $N^{D}$ and $N^{S}$ represent the normal force carried by the two asperities: $N=N^{D}+N^{S}$. The corresponding local friction coefficients are $\mu^{D}$ and $\mu^{S}$, respectively. In the second scenario, the difference in the normal forces carried by the two asperities is smaller (Fig. 4(d)). This results in both contact asperities being stuck at the critical state of the interface. Now, the failure of the weaker asperity will trigger the slip of the stronger asperity due to its excess friction force redistribution. The asperity with lower normal force has reached a coefficient of $\mu^{S}$, while the contact asperity with higher normal force experiences a local friction coefficient of $\mu^{0}$ that is smaller than $\mu^{S}$. Note that contrarily to $\mu^{D}$ and $\mu^{S}$, $\mu^{0}$ is not a fixed value related to the nanoscale friction experiment: it is the instantaneous friction coefficient experienced by a stuck asperity that is loaded sub-critically. The calculated interfacial friction coefficient is then $\mu=(\mu^{0}N^{0}+\mu^{S}N^{S})/N$, where $N^{0}$ and $N^{S}$ represent the normal force on the two asperities, satisfying $N=N^{0}+N^{S}$ and $N^{0}>N^{S}$. In both scenarios the static friction coefficient is lowered by the addition of a second asperity, either because an asperity has already entered slip when the overall friction force is maximal (scenario 1) or because the stronger asperity is “subcritical”; its slip is activated by failure of the weaker asperity (scenario 2).
@@
We can therefore define three types of contact asperities at the critical state of multi-asperity interfaces: critical contact asperities, defined as those that experience the local static friction coefficient of $\mu^{S}$; pre-sliding contact asperities, defined as those that have already slipped and therefore show the dynamic friction coefficient of $\mu^{D}$; and subcritical contact asperities, defined as those with a local friction coefficient of $\mu^{0}$ that is smaller than $\mu^{S}$. For a realistic multi-asperity interface, many asperities might correspond to each of the three asperity types. An example is shown in Fig. 4(e), with the distribution of local friction coefficients for an interface at the critical state under a normal load of 10 mN. The three types of asperities (critical, pre-sliding, and subcritical) are respectively represented in green, yellow, and brown. The corresponding contact map from the BIM computation is shown in Fig. 4(f), with the contact being colored following the same scheme.
@@
The slip of the subcritical asperities is activated by redistribution of excess friction force released when the critical asperities slip. This instability in our model defines the critical interface state with maximal applied tangential load. Since both $\mu^{D}$ and $\mu^{0}$ are smaller than $\mu^{S}$, the interfacial friction coefficient at the critical interfacial state will, as expected, be smaller as compared to that in a single asperity interface ($\mu<\mu^{S}$). Consequently, it is the pre-sliding and subcritical contact asperities that dominate the reduction of the macroscopic static friction coefficient at the multi-asperity interfaces with respect to its value for a single asperity interface, leading to a smaller friction drop with increasing normal load. In other words, the critical asperities govern the maximum static friction coefficient. To quantify the relative contribution of critical, pre-sliding, and subcritical contact asperities to the overall static friction coefficient, we plot the fraction of the total normal load (respectively friction force) supported by different types of contact asperities under various total normal loads in solid (respectively dashed) lines in Fig. 4(g). As the total normal load increases, more contact asperities form at the contact interface. The fraction of pre-sliding and subcritical contact asperities gradually increases from 0 at single asperity contact interface to an approximately constant value at multi-asperity contact interface, which suggests a greater reduction of static friction coefficient and thus results in a smaller macroscopic static friction coefficient. In contrast, the fraction of critical contact asperities decreases from 1 to an almost negligible level. This decreasing fraction of critical contact asperities with increasing normal load indicates a smaller contribution of critical contact asperities to the static friction coefficient, which matches well with the observed friction drop in Fig. 2. We expect the range of load over which the fraction of critical asperities is replaced by pre-sliding and subcritical asperities to depend on the bulk parameters and the geometric details of the surfaces, for example, their roughness.

### v2

# Supplemental_material

@@ Abstract
The key parameter for describing frictional strength at the onset of sliding is the static friction coefficient. Yet, how the static friction coefficient emerges at the macroscale from contacting asperities at the microscale is still an open problem. Here, we present friction experiments in which the normal load was varied over more than three orders of magnitude, so that a transition from a single asperity contact at low loads to multi-asperity contacts at high loads was achieved. We find a remarkable reduction in the friction drop (the ratio of the static friction force to the dynamic friction force) with increasing normal load. Using a simple stick-slip transition model we identify the presence of pre-sliding and subcritical contact points as the cause of smaller static friction coefficient at increased normal loads. Our measurements and model bridge the gap between friction behavior commonly observed in atomic force microscopy (AFM) experiments at microscopic forces, and industrially relevant multi-asperity contact interfaces loaded with macroscopic forces.
@@
In this Letter, we experimentally demonstrate that the friction drop (the ratio of static to dynamic friction coefficient) decreases as the interface grows from single contact to multi-asperity contact by increasing the normal load applied to a silicon-on-silicon contact. The observed decreasing trend of friction drop with increasing normal load can be captured by a simple stick-slip transition model, analogous to a fiber bundle model, based on the local normal contact stress map obtained from boundary integral contact calculations without adjustable parameters. We find that three types of contact points can be identified when the interface is subjected to the maximal tangential force it can withstand. We describe these three types of contact elements as “critical”, “pre-sliding” and “subcritical”. While single-asperity interfaces are always critical at the onset of slip, we show that the emergence of pre-sliding and subcritical contact asperities at multi-asperity interfaces causes the friction drop to decrease.
@@
Figure 1: Experimental system. (a) A 3 mm-diameter silicon ball is clamped to the geometry of a rheometer and brought into contact with a homemade Si cantilever. By rotating the geometry at a constant angular velocity, $\omega$, a sliding speed of $\omega r$ ($r=10\text{ mm}$) is imposed at the Si-on-Si contact interface. The corresponding friction force ($F_{f}$) and the normal force ($N$) can be simultaneously measured by the rheometer and the balance. The ratio of friction force and normal load gives the coefficient of friction, $\mu={F_{f}}/{N}$. (b) An example of the slide-hold-slide experiments protocol with a normal load of 24 mN and a hold time of 100 s. $F_{S}$ and $F_{D}$ represent the maximum static friction force and the average dynamic friction force after the sliding is reinitiated, respectively.
@@
We study silicon-on-silicon friction using a rheometer (DSR 502, Anton Paar) in combination with an analytical balance (Quintix, Sartorius) inside a humidity-controlled chamber, see Fig. 1(a). To prepare highly hydroxylated silicon surfaces [16, 15], the silicon balls and the silicon cantilevers were treated by oxygen plasma cleaning inside a reactor (Zepto One, Diener Electronic) [23]. In the friction experiments, the normal loads were precisely varied from 0.1 to 100 mN by using silicon cantilevers of different normal stiffness (see more details in Appendix A ). The imposed sliding speed was fixed at 0.3  µm/s for all the measurements. To reset contact aging [34, 17, 16], a slide-hold-slide (SHS) measurement procedure was employed (see Fig. 1(b)): we used the rheometer to impose a stationary state with zero-torque (or zero-friction) after the initial sliding to avoid impacting the aging process with shear force [35, 4, 17], followed by reinitiating the sliding in the same direction. The static friction force $F_{S}$ and dynamic friction force $F_{D}$ when sliding was restarted, were recorded to calculate the corresponding static friction coefficient, $\mu_{S}$, the dynamic friction coefficient, $\mu_{D}$, and the friction drop, $\mu_{S}/\mu_{D}$, which equals to $F_{S}/F_{D}$. All sliding strokes were measured on previously untouched areas of the silicon wafer and kept short at 10 µm to minimize the influence of wear (Fig. S2).
@@
Both the static and the dynamic friction coefficients, defined as the ratio of friction force to normal force, are widely assumed to be independent of the normal load [36, 9, 37]. Surprisingly, we observe that the friction drop gradually decreases with increasing normal load as shown with black data points in Fig. 2. Notably, the nanoscale single-asperity friction drop (leftmost black circle in Fig. 2), reproduced from previous AFM experiments (inset of Fig. 2) [15], aligns well with our observed qualitative decreasing trend of friction drop with increasing normal load.
@@
Figure 2: The friction drop as a function of normal load, measured from experiments (data points) and predicted by our model (shaded area), with black and red corresponding to a hold time of 100 s and 10 s, respectively (the static and dynamic friction coefficients are reported in Fig. S1.). Each symbol stands for a different Si cantilever with varying normal stiffness, and each data point represents the average of three measurements. The leftmost red and black solid circle data points, measured from AFM experiments, are extracted from Fig. S5 in Ref. [15], with the evolution of the nanoscale friction coefficient being reproduced in the inset. The vertical lines indicate the normal load corresponding to the normal stress distribution shown in Fig. 3.
@@
To provide insight into the normal load dependence of the friction drop, we propose a simple stick-slip transition model. In the model, we discretize the interface with square elements (pixels) such that the macroscopic friction force is the sum of all the microscopic interfacial shear forces generated by these elements. We also assume that the element-scale frictional behavior is characterized by an instant drop in friction upon slip, the magnitude of which matches that observed in the AFM friction experiments shown in the inset of Fig. 2. In other words, the local ratio of shear stress to normal stress (local stress ratio) drops at the onset of local slip, in accordance with the static and dynamic friction coefficient ($\mu^{S}$ and $\mu^{D}$, respectively) observed in the AFM experiments. The interface is a collection of microscopic contact patches of varying size (see Fig. S4). In our model, the shear force applied to an individual contact patch scales linearly with the number of elements involved in that patch. This means that the applied shear force is distributed homogeneously over the real contact area before any element starts to slip. The excess shear force generated when an element enters slip is uniformly redistributed over the interfacial contact area (elements) that has not undergone slip yet. We also assume that there is no evolution of the area and geometry of the contact patches during sliding, as the substrate wafer is much smoother than the silicon ball [23].
@@
To estimate the local normal stress at the interface, we conduct elastic contact calculations using an open-source implementation of the boundary integral method (BIM) [38]. In the contact calculation, the elastically stored energy of the contact system is minimized under the constraint that the resulting mean contact pressure balances the applied average pressure over the area of the input topography (see more details in Appendix B Our use of elastic contact is reasonable here, since there is mainly elastic deformation at the contact interface (see Fig. 3) [30, 23, 39]. Using the calculated local normal stress distribution and the assumed relations between local shear stress and local normal stress described above, the model predicts how the total tangential force, which is externally imposed, is distributed over the contact area and how this distribution of local shear stress evolves as the applied tangential force is increased. In our model, at the level of a single element, the maximal (critical) ratio of shear stress to normal stress is constant; in line with the load-controlled friction framework [40, 41, 42]. It is the normal stress distribution that completely determines the macroscopic frictional behavior. This enables us to predict the frictional response of the same interface at various normal loads, from which we can compute the macroscopic static friction coefficient and friction drop.
@@
In other words, our model is analogous to a fiber bundle model with equal load sharing [43], where the failure of a bundle of fibers corresponds to the failure of contact elements, with their strength governed by the local normal stress. While sliding heterogeneity can originate from elastic interaction among contact patches [11], we assume that these interactions during the onset of the sliding process are negligible in our system. Indeed, the nanometer-scale contact deformation ($\sim$ 5 nm at a normal load of 10 mN based on BIM contact calculation as mentioned above) is considerably smaller than the micrometer-scale distance between contact patches. Therefore, it is reasonable to ignore the role of asperity interactions in the distribution of the shear forces. However, this assumption may break down at increasing normal loads when the distances between contact patches shrink (see Fig. S4).
@@
To quantify ball-to-ball topography variations, we use three different topographies measured on silicon balls that were used in the experiments (see Fig. S3). For each topography and normal load, the model predicts the friction drop. The gray shaded area in Fig. 2 corresponds to the range of predicted friction drop values based on the three topographies, and the nanoscale static and dynamic friction coefficients measured at a hold time of 100 s in AFM experiments. The stick-slip transition model qualitatively reproduces the decrease of the friction drop with increasing normal load and quantitatively matches the experiments without any fitting parameter.
@@
To evaluate the robustness of our model against variations in the system characteristics, we conduct additional experiments with a hold time of 10 s instead of 100 s, and report the corresponding friction drop with red data points in Fig. 2. A decrease of the friction drop with increasing normal load is observed, similar to the experiments for a hold time of 100 s. Furthermore, it is noteworthy that the measured friction drop for a hold time of 100 s is larger than that for a hold time of 10 s under the same normal force. The larger static friction coefficient, corresponding to a greater friction drop, is expected: a longer hold time during the SHS procedure can enhance the contact strength by forming more interfacial siloxane bonds [15, 23], which results in a higher static friction coefficient. Moreover, our stick-slip transition model (red-shaded area in Fig. 2) with the nanoscale static and dynamic friction coefficients measured at a hold time of 10 s from previous AFM experiments [15] (leftmost red circle in Fig. 2, see also Fig. S5), still captures successfully the evolution of friction drop with the normal force for the experiments with 10 s hold time. One should note that there is still a gap between the minimum normal load used in the simulation and AFM experiments in Fig. 2 (marked by black and red horizontal dot lines), which will be discussed later.
@@
To further elucidate the relation between the normal load, the local normal stress distribution, and the macroscopic frictional response, we plot the probability density function, $p(\sigma)$, of the normal stress at the interface, obtained from elastic contact calculations with various normal loads in Fig. 3. With increasing normal load, the probability density function of the local normal stress evolves from being discrete (the black curve, 0.1 mN) to continuous (the yellow curve, 10mN). This directly relates to the transition from single to multi-asperity contact interfaces (see Fig. S4). At lowest normal loads, only a few contact patches are formed, and the macroscopic behavior may be dominated by those patches. At larger normal loads, however, there is a more continuous distribution of normal (and thus shear) stresses at the interface. The normal loads corresponding to the distribution shown in Fig. 3 are indicated with vertical lines of the same color in Fig. 2.
@@
Figure 4: Local stick-slip transition at the critical state of the interface from numerical computation. (a) Macroscopic friction coefficient versus displacement curve at a normal load of 0.1 mN. The black circle corresponds to the critical interfacial state where no more additional tangential load can be added. (b) to (e) Distribution of local stress ratio for a single-element interface (b), two-element interface (c)(d), and multi-element interface under a normal load of 10 mN (e), respectively. The brown, yellow, and green correspond to critical, pre-sliding, and subcritical elements with local stress ratios of $\mu^{D}$, $\mu^{S}$, and $\mu^{0}$, respectively. $\mu^{0}$ is smaller than $\mu^{S}$. $n^{\mu}$ and $n^{tot}$ represent the number of elements under a local stress ratio of $\mu$ ($\mu=\mu^{S}$, $\mu^{D}$, or $\mu^{0}$) and the total number of elements, respectively. Movies of the evolution of the distribution of local stress ratios with increasing imposed displacement are available in Appendix C. (f) The state of the contacting elements at the critical interfacial state for a normal load of 10 mN. (g) The fraction of the total normal force (solid lines) and the total friction force (dashed lines) generated by different types of contact elements in the interfaces under various normal loads. $N^{\mu}$ and $N$ represent the shared normal load on those contact elements under a local stress ratio of $\mu$ ($\mu=\mu^{S}$, $\mu^{D}$, or $\mu^{0}$) and the total applied normal load, respectively. $F_{f}^{\mu}$ and $F_{f}$ represent the same quantity but for the shear load.
@@
To better understand how the differences in normal stress distribution impact the onset of sliding, we focus on the critical state of the interface, where the peak macroscopic friction force is reached under an imposed tangential lateral displacement of $\delta_{\text{peak}}$ (see a typical simulated friction-displacement curve in Fig. 4(a)). We plot the distribution of local stress ratios across all the contact elements at the critical state of the interface in Fig. 4(b)-(e). For an idealized single element contact (Fig. 4(b)), the element will reach the maximum local stress ratio, $\mu^{S}$, resulting in an interfacial friction coefficient equal to the local stress ratio ($\mu=\mu^{S}$). In this limiting case of a single asperity contact, one recovers the frictional behavior as measured from AFM. In our model, this means that below a normal load for which the real contact area is of the scale of the input AFM topography resolution (21 nm), the macroscopic friction coefficient is not dependent on the normal load. Such an extreme case is marked by black and red horizontal dot lines in Fig. 2.
@@
The situation becomes more complex when considering multi-asperity contact interfaces, as each element potentially experiences different local normal stresses (and thus different local stress ratios at the critical interfacial state). We stress that, within our model, any distribution of local stress ratios (between 00 and $\mu^{S}$) at a multi-contact interface can only result in a decrease in the friction drop as compared to the friction drop at single-asperity level. Hereafter we investigate the details of the critical interfacial state for multi-asperity interfaces and its influence on the friction drop, starting with the idealized case of two contacting elements.
@@
For an interface composed of two contacting elements, each carrying a different normal stress, two distinct scenarios arise at the critical interfacial state corresponding to maximal (global) static friction force. In the first scenario, the normal stress carried by each element varies significantly (Fig. 4(c)) such that the weaker element, carrying a smaller normal stress, has already slipped and experiences local stress ratio $\mu^{D}$, while the stronger element with the higher normal stress remains stuck with a local stress ratio of $\mu^{S}$. We can calculate the interfacial friction coefficient as $\mu=(\mu^{D}N^{D}+\mu^{S}N^{S})/N$, with $N^{D}$ and $N^{S}$ the normal forces carried by the two elements: $N=N^{D}+N^{S}$. In the second scenario, the difference in the normal stresses carried by the two elements is smaller (Fig. 4(d)). This results in both contact elements being stuck at the critical state of the interface. Now, the failure of the weaker element will trigger the slip of the stronger element due to its excess friction force redistribution. This is reminiscent of the asperity model in geophysics [44], in which the failure of one asperity may cause an avalanche of asperity failure. The element with lower normal stress has reached a local stress ratio of $\mu^{S}$, while the contact element with higher normal stress experiences a local stress ratio of $\mu^{0}$ that is smaller than $\mu^{S}$. Contrarily to $\mu^{D}$ and $\mu^{S}$, $\mu^{0}$ is not a fixed value related to the nanoscale friction experiment: it is the instantaneous local stress ratio experienced by a stuck element that is loaded sub-critically. The collective friction coefficient is then $\mu=(\mu^{0}N^{0}+\mu^{S}N^{S})/N$, where $N^{0}$ and $N^{S}$ represent the normal force on the two elements, satisfying $N=N^{0}+N^{S}$ and $N^{0}>N^{S}$. In both scenarios the friction drop is lowered by the addition of a second element, either because an element has already entered slip when the overall friction force is maximal (scenario 1) or because the stronger element is “subcritical”; its slip is activated by failure of the weaker element (scenario 2).
@@
We can therefore define three types of contact elements at the critical state of multi-asperity interfaces: critical contact elements, experiencing the maximal local stress ratio $\mu^{S}$; pre-sliding contact elements, that have already slipped and therefore show the dynamic stress ratio $\mu^{D}$; and subcritical contact elements, with a local stress ratio of $\mu^{0}$ smaller than $\mu^{S}$. For a realistic multi-asperity interface, many elements might correspond to each of the three element types. An example is shown in Fig. 4(e), with the distribution of local stress ratios for an interface at the critical state under a normal load of 10 mN. The three types of elements (critical, pre-sliding, and subcritical) are respectively represented in green, yellow, and brown. The corresponding contact map from the BIM computation is shown in Fig. 4(f), with the contact being colored following the same scheme.
@@
The slip of the subcritical elements is activated by the redistribution of excess friction force released when critical elements slip. This instability in our model defines the critical interface state with maximal applied tangential load. Since both $\mu^{D}$ and $\mu^{0}$ are smaller than $\mu^{S}$, the collective friction coefficient at the critical interfacial state will, as expected, be smaller as compared to that in a single element interface ($\mu<\mu^{S}$). Consequently, it is the pre-sliding and subcritical contact elements that dominate the reduction of the collective static friction coefficient at the multi-asperity interfaces with respect to its value for a single element interface, leading to a smaller friction drop with increasing normal load. To quantify the relative contribution of critical, pre-sliding, and subcritical contact elements to the overall static friction coefficient, we plot the fraction of the total normal load (respectively friction force) supported by different types of contact elements under various total normal loads in solid (respectively dashed) lines in Fig. 4(g). As the total normal load increases, more contact elements form at the contact interface. The fraction of pre-sliding and subcritical contact elements gradually increases from 0 at single element contact interfaces to an approximately constant value at multi-asperity contact interfaces, which suggests that the friction drop decreases with increasing normal force. In contrast, the fraction of critical contact elements decreases from 1 to an almost negligible level. This decreasing fraction of critical contact elements with increasing normal load indicates a smaller contribution of critical contact elements to the static friction coefficient, which matches well with the observed friction drop in Fig. 2. In the low normal load limit, the evolution of friction drop is given by the intrinsic, nano-scale friction drop which can for example be controlled by interfacial bonding dynamics [24]. At low normal loads, the critical asperities govern the static friction coefficient, which is analogous to the dominant role of nucleation regions as identified in previous work [45, 46]. The specific normal load range over which the friction drop decreases, is controlled by the elastic properties and surface roughness in our model, with the relation between friction drop and normal load shifting to higher normal loads for higher roughness or stiffer materials as long as the contact mechanics remains elastic. For example, MEMS devices typically involve lower roughness and lower normal force compared to our experimental conditions, and may therefore display similar friction drop behavior. At larger interfaces, however, additional complexity may arise [27, 47]. The assumption of independent contact elements in our model may break down at larger loads, where large contact patches lead to increased interfacial stiffness and shear loading [48], resulting in an even more heterogeneous distribution of shear stress, which could further reduce the predicted friction drop.